# On Computability of Computable Problems

### v1

@@ 6 Mathematical Formulation of Stipulation Machine Interactions
Stipulation machine and Turing machine are realizable objects, and if they interact with each other through eight mutually binding rules (postulates) then their interactions can be described through $f(Z,T)$ transform.We apply $f(Z,T)$ transform only on computable problems, and transformed problems can be computable (case I and II) or uncomputable (case IV), and sometime the computability of the transformed problem is highly sensitive toward choice of computational scheme(case III).
7 Discussion
This paper presents a new realizable computational scheme and through this scheme we transform Turing computable problems into new type of problems.The transformed problems provide a window to analyse the computational characteristics of different variants of Turing machine at infinitesimal scale, and to enlist those variants of Turing machine according to their computational characteristics.This transform has some applied aspects as well.The $f(Z,T)$ transform describes interactions between Turing machine (or Turing complete models of computation) and Stipulation machine that take place under eight postulates.
There is a problem $P$ and it is computable on input $\psi$.We construct $f(Z,T)$ transform that transforms computable problem $P(\psi)$ into new problem $P(t)(\psi)$.Then we analyse the computability of transformed problem $P(t)(\psi)$ through different variants of Turing machines using $f_{n}(Z,T)$ normalization line.If we turn $P(\psi)$ into $P(t)(\psi)$ and $10\leq f_{n}(P,T)<$ 1 then $P(t)(\psi)$ is also computable just like $P(t)(\psi)$, but if $f_{n}(P,T)>1$ then $P(t)(\psi)$ is not computable problem.If $f_{n}(P,T)=1$ then $P(t)(\psi)$ exist at point of inflection, and this point provides a very important opportunity to investigate computational characteristics of different variants of Turing machines at infinitesimal scale.If we turn $P(\psi)$ into $P(t)(\psi)$ and $f_{n}(P,T)=0$ then $P(t)(\psi)$ exists at point of convergence.If $P(t)(\psi)$ exists at point of convergence then $T(k,w)=T(1,1)$ for every finite $k$ and $w$; i.e., finite number of additional tapes or additional alphabets do not provide additional computational power to a Turing machine.We apply $f(Z,T)$ transform on a computable problem $P(\psi)$ and we get transformed problem $P(t)(\psi)$.Following two equations represent the characteristics of this $f(Z,T)$ transformed problem.
$f(P,T)=\frac{N_{P}T_{t}}{1-2^{-\left(\nicefrac{{T_{P}}}{{N_{P}T_{t}}}\right)}}$ (37)
$f_{n}(P,T)=\frac{N_{P}T_{t}}{T_{P}}$ (38)
There are two expressions in (37 and 38) that represent time, i.e., $T_{t}$ and $T_{P}$.The time $T_{t}$ is characteristic of Turing machine and time $T_{P}$ is characteristic of Stipulation machine.If $(\nicefrac{{T_{t}}}{{T_{P}}})$ such that $f_{n}(P,T)>1$ then $P(t)(\psi)$ is uncomputable, and if $(\nicefrac{{T_{t}}}{{T_{P}}})$ such that $f_{n}(P,T)<1$ then $P(t)(\psi)$ is also computable, like $P(\psi)$. As stated earlier, if $f_{n}(P,T)=0$ then finite number of additional tapes or additional alphabet do not provide any additional computational power to Turing machines, and $P(t)(\psi)$ is computable by every Turing machine defined over finite number of tapes and finite set of alphabet. Therefore, it can be stated that established results in computability theory are special case of $f(Z,T)$ transform where $f_{n}(P,T)=0$. If problem $P(\psi)$ is a computable then problem $P(t)(\psi)$ with $T_{P}=\infty$ is also computable.Therefore, it can also be stated that $P(\psi)=P(t)(\psi)$ when $T_{P}=\infty$, and consequently, $T(1,1)=T(k,w)$, over finite number of $k$ and $w$.We can say that every $P(\psi)$ is in fact $P(t)(\psi)$ whose input never changes while it is being computed.The $f(Z,T)$ transform, through (37 and 38), establishes a notion of relativistic computability within computable problems.In this relative computability, if a problem is computable for a given inputs and inputs are changing with time, if the rate of change of inputs is sufficiently high for given model of computation, then the computable problem turns into uncomputable problem for the given model of computation. Suppose, there is a task $G$ and it is computable if and only if all given information (input, instructions, etc.) are taken into account.The task $G$ gets its information $\eta$ through an interface.We apply $f(Z,T)$ transform on $G(\eta)$ and we get (39 and 40).
$f(G,T)=\frac{N_{G}T_{t}}{1-2^{-\left(\nicefrac{{T_{G}}}{{N_{G}T_{t}}}\right)}}$ (39)
$f_{n}(G,T)=\frac{N_{G}T_{t}}{T_{G}}$ (40)
The $T_{G}$ in (39 and 40) represents time after which input information $\eta$ varies.The task $G(t)(\eta)$ with fixed input information has $f_{n}(G,T)=0$ because $T_{G}=\infty$ and it is computable task.If we start changing the input and $T_{t}\ll T_{G}$ such that $f_{n}(G,T)<1$ then $G(t)(\eta)$ remains a computable task, and if we start changing input in a way that $f_{n}(G,T)>1$ then $G(t)(\eta)$ turns into uncomputable task.If the task $G(t)(\eta)$ represents a fixed size phenomenon that evolves with time (as in 13 and 27 ) then $G(t)(\eta)$ is computable only if $f_{n}(G,T)<1$.Stipulation and Turing machines can be designed to replicate the phenomena that are based on $f(Z,T)$ transform.The phenomena that are based on $f(Z,T)$ Transform are realizable, and consequences of those phenomena are observable. Future work will extend the results of $f(Z,T)$ Transform toward complexity theory, and it will also present the applied aspects of this Transform.

### v3

@@ 5 Mathematical Formulations of $f(Z,T)$ Transform
We apply $f(Z,T)$ transform only on computable problems, and transformed problems can be computable (case I and II) or uncomputable (case IV), and sometime the computability of the transformed problem is highly sensitive toward choice of computational scheme(case III).If $T_{P}=\nicefrac{{1}}{{F_{P}}}$ and $T_{t}=\nicefrac{{1}}{{F_{t}}}$ then we can represent normalisation of transformed problem through frequencies in following way.
$f_{n}(P,F)=\frac{N_{P}F_{P}}{F_{t}}$
There are two frequencies in $f_{n}(P,F)$, i.e., $F_{t}$ and $F_{P}$.The frequency $F_{t}$ is related to clock speed of Turing machine and $F_{P}$ describes the frequency of recurring inputs.If frequency of recurring inputs is higher than certain threshold frequency then computable problem turns into uncomputable problem.It is helpful to describe normalization through frequencies if transformed problems involve measuring some frequencies.
6 Summary
In this paper we develop a transform technique and apply it on some decidable language, and show that cardinality of transformed languages is uncountable.There are countably infinite Turing machines but uncountable number of transformed languages, and this difference in cardinality implies the existence of transformed problems that are not computable.Furthermore, we present a realizable computational scheme that can demonstrate $f(Z,T)$ transform, and this scheme consists of Turing machine and Stipulation machine.Through $f(Z,T)$ transform, quite unexpectedly, we will be able to formalize some important natural phenomena.We will present experimental validation of this transform in future work, and those experiments will be designed and executed in such a way that it shall be easy to reproduce them.